Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 21:36 Tuesday 16 March 2021 (16272) Print this report
Tightening of the wire clamp screws.

With Hirata-san

See pictures in album PR3 Remedying Work.

The goal is to tighten the wire clamps at the sides of the IM with the hope of solving the pitch instability problem. 

The procedure we followed was as follows:

  1. Lock the BF side to the security structure.
  2. Raise the IRM lock screws brackets to the height of the IRM top panel.
  3. Lock the IRM from the top panel trying to keep it at the current free-hanging position.
  4. Lock the IM at the current free-hanging position.
  5. Lock the RM to the security structure.
  6. Lock the mirror to the RM.
  7. Disconect  the H2 and H3 OSEM cables and put them on the top panel.
  8. Remove H2 and H3 OSEM flags.
  9. Remove the side panels.
  10. Inspect the clamps visually.
  11. Tighten the weight holding clamps M6 screws to 10 Nm.
  12. Tighten the locating clamps M5 screws to 2 Nm.


  • Today we reached step 11. Tomorrow we will do step 12 and reassemble the IRM.
  • The screws of the locating clamps for the piano wire holding the mirror are not accesible. See this picture
  • Taking advantage that we removed IRM panels, we took pictures of the ballast masses on top of the IM, just in case we need to know their mass values for another task. See the pictures of the day in the album.
Comments to this report:
fabian.arellano - 16:56 Wednesday 17 March 2021 (16277) Print this report

With Hirata-san and Washimi-san.

See pictures in album PR3 Remedying Work.

Summary: after tightening the wire clamps we tested the pitch stability using the Trump method. However, the situation did not change. Pitch still changes a lot when the suspension is shaken.

The work sequence was this:

  1. We tightened the positioning clamp screws for the RM wires to 2 Nm.
  2. We put back the side IRM panels with the OSEMs.
  3. We put the OSEM flags in place.
  4. We released the payload and thew BF side.
  5. We inspected the horizonal OSEM flags visually and saw there was a large negative pitch. H1 and H3 flags were touching the OSEM bodies.
  6. We corrected the pitch with the picomotors and we also modified the roll a little using the readout from V2 and V3 OSEMs, which were within range. The IM became free from touching the IRM.
  7. Transfer function measurements in L, Y and R suggested the IM was free. The coherence was low a low frequencies likely because the IM was still moving a lot due to us working with it.
  8. We set up an oplev at the back of the mirror.
  9. We applied the Trump procedure.

Upon the procedure, according to the oplev, the mirror moved in pitch approximately by -5.8 mrad, whereas, according to OSEMs V2 and V3, the IM moved by -1.04 mrad. See the details of both calculations in this notebook page.

Note: although it was clarified in the commented entry, it's worth writing here that the screws of the locating clamps for the piano wire holding the mirror are not accesible. See this picture

Search Help